[Brahms-l] Shift report 20040216 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 16 2004 - 23:44:06 EST
------------------------------------------------------------
Date: Mon Feb 16 23:44:06 2004, Shift: 16:00-24:00
Supervisor: IGB, with: ZIS
Log book#15 pages 60 - 60

Shift Summary:
------------------------------------------------------------
1. Beam has been on from the entry of shift. 
 The background was high, and we voted dump at 16:32.
 New beam started arround 17:18 until 19:30 when that has 
 been not steered.
 Few attemps, but STAR requested access in IR.     

2. Runs 10395,10396,10398,10400 trg2(347K), tr3(500k)

3. No detector problems
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 16 23:44:23 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 16 2004 - 23:44:43 EST